Two particles with charges +5e and -5e are initially very far apart (effectively an infinite distance apart). They are then fixed at positions that are 7.71 x 10-10 m apart. What is EPEfinal - EPEinitial, which is the change in the electric potential energy?
Answer #1

Since the two charges are far apart

the electric potential energy is given by

EPE = \frac{kq_1q_2}{r}

so initially the distance r = infinite

EPE_i = \frac{kq_1q_2}{r} = 0

Now final position r = 7.70*10^{-10}m

EPE_f = \frac{9*10^9*5e*(-5e)}{7.7*10^{-10}}

EPE_f = -7.5*10^{-18}J

Now we have

EPE_f - EPE_i = -7.5*10^{-18}J